OpenStudy (anonymous):

Bill comments that the weight of his textbooks is only 7 pounds more than ½ the weight of Maria’s textbooks. The total weight of Bill and Maria’s textbooks is 25 pounds. How could this be written as an algebraic expression to determine the weight of Maria’s textbooks?

OpenStudy (calculusxy):

7+(x*1/2)=25 x stands for the number of pages in Maria's textbook

OpenStudy (anonymous):

That is what I thought was the answer but it keeps saying the answer is 3/2 M +7 =25.

OpenStudy (anonymous):

Bill's weight of his textbook = 1/2 (maria's textbooks' weight) + 7 maria's textbooks' weight + Bill's weight of his textbook is = 25 pounds

OpenStudy (anonymous):

Do substitution from the first equation to the second one

OpenStudy (anonymous):

So...1/2M+7=25

OpenStudy (calculusxy):

This is the same thing as (1/2*M)+7=25

OpenStudy (anonymous):

So my answer was right and the answer on the practice test is wrong.

OpenStudy (anonymous):

well what is the answer on the practice test?

OpenStudy (anonymous):

3/2M +7 =25

OpenStudy (calculusxy):

@staldk3 Did the question say such thing as 3/2 of the pages at Maria's textbook.

OpenStudy (anonymous):

that is the actual answer because (1/2*M)+7 is only bill's textbooks' weight and 25 = both bill's and maria's

OpenStudy (anonymous):

Bill comments that the weight of his textbooks is only 7 pounds more than ½ the weight of Maria’s textbooks. The total weight of Bill and Maria’s textbooks is 25 pounds. How could this be written as an algebraic expression to determine the weight of Maria’s textbooks?

OpenStudy (anonymous):

(1/2*M)+7 + M = 25 3/2M + 7 = 25

OpenStudy (calculusxy):

But iisn't 3/2 stands for 1 1/2

OpenStudy (anonymous):

it can be both

OpenStudy (calculusxy):

Okay then

OpenStudy (anonymous):

I don't really understand how you get 3/2.

OpenStudy (calculusxy):

Me neither

OpenStudy (anonymous):

1/2 M + M = ?

OpenStudy (anonymous):

But M is just a letter...

OpenStudy (anonymous):

its a variable so in this case you would collect like terms so that it would be (1/2 + 1) M = 3/2 M

OpenStudy (anonymous):

So the M =1?

OpenStudy (anonymous):

no M does not have a value atm

OpenStudy (anonymous):

Then how did you get the 1?

OpenStudy (anonymous):

1M is like the same as M

OpenStudy (anonymous):

Okay. That makes sense. It was just an imaginary 1.

OpenStudy (anonymous):

Thanks!